# R&D Networks under Heterogeneous Firm Productivities

### v1

@@ Appendix A Proofs
Start from a fully connected network $\mathbf{G}$ such that $\mathbf{G}_{ij}=1$ for all $i,j\in\mathcal{N}$ with $i\neq j$. For this network to be pairwise stable, there must not be any profitable linking deviations. In the complete case, any deviation involves deleting a link. Without loss of generality, consider deleting $G_{1,2}$ by setting it equal to 0.
First, solve the system of equations in (7). Since $\forall i,j\in\mathcal{N}:G_{ij}=1$, we have
$\left[\frac{1}{\theta_{i}}\frac{(n+1)^{2}\phi}{n-d_{i}(\mathbf{G})}-(n+1)\theta_{i}\right]e_{i}=(\alpha-\bar{c})+\sum_{j=1}^{n}\theta_{j}e_{j}\Big[(n+1)-(1+\delta d_{j}(\mathbf{G}))\Big].$
Using the fact that in a complete network $d_{i}(\mathbf{G}^{C})=n-1,\forall i\in\mathcal{N}$, we obtain
@@
${e}^{*}_{i}(\mathbf{G}^{C})=(\alpha-\bar{c})\cdot\frac{\theta_{i}}{(n+1)^{2}\;\phi-\sum_{k\in\mathcal{N}}\theta_{k}^{2}},\quad\forall i\in\mathcal{N}.$ (15)
By deviating from the complete network and removing the link between agents $i$ and $j$, we obtain
${e}^{*}_{i}\big(\mathbf{G}^{C}_{-ij}\big)=(\alpha-\bar{c})\cdot\frac{\lambda}{\big[1-(1+\Delta)\,Q\big]\cdot\big[b_{i}\,\lambda-(n-1)\,\theta_{j}\big]},$ (16)
where
$\begin{gathered}\lambda=\frac{\theta_{i}}{\theta_{j}}\cdot\frac{(n+1)^{2}\phi-2\,\theta_{j}^{2}}{(n+1)^{2}\phi-2\,\theta_{i}^{2}},\qquad\Delta=\frac{2\,(\lambda\,\theta_{i}+\theta_{j})}{b_{i}\,\lambda-(n-1)\theta_{j}},\\ Q=\sum_{k\in\mathcal{N}/\{i,j\}}\frac{\theta_{k}^{2}}{(n+1)^{2}\,\phi},\qquad b_{i}=\frac{(n+1)^{2}\,\phi}{2\,\theta_{i}}-2\,\theta_{i}.\end{gathered}$
Using the equilibrium profit in Equation (9), the profit for firm $i$ under full connectivity is
@@
${\pi}^{*}_{i}\big(\mathbf{G}^{C}\big)=\left[\frac{\phi}{\theta_{i}^{2}}\,(n+1)^{2}-1\right]\,\phi\,\left((\alpha-\bar{c})\cdot\frac{\theta_{i}}{(n+1)^{2}\,\phi-\sum_{k\in\mathcal{N}}\theta_{k}^{2}}\right)^{2},$ (17)
while after deleting the link $(i,j)$ it becomes
${\pi}^{*}_{i}\big(\mathbf{G}^{C}_{-ij}\big)=\left[\frac{\phi}{\theta_{i}^{2}}\left(\frac{n+1}{2}\right)^{2}-1\right]\,\phi\,\left((\alpha-\bar{c})\cdot\frac{\lambda}{\big[1-(1+\Delta)\,Q\big]\cdot\big[b_{i}\,\lambda-(n-1)\,\theta_{j}\big]}\right)^{2}.$ (18)
Assume that $\theta_{i}$ represents the productivity of the most productive firm, i.e. $\theta_{i}=1$, and that $\theta_{j}$ is the least productive firm. In the limit $\theta_{j}\to 1$, all firms are homogeneous with $\theta_{\min}=\theta_{\max}$. This corresponds to the case in goyal2001r, where it is shown that deviation from a complete network is not beneficial:
$\lim_{\theta_{j}\to 1}{\pi}^{*}_{i}\big(\mathbf{G}^{C}\big)>\lim_{\theta_{j}\to 1}{\pi}^{*}_{i}\big(\mathbf{G}^{C}_{-ij}\big).$
Now consider the opposite case $\theta_{j}\to 0$. We obtain
$\lim_{\theta_{j}\to 0}{\pi}^{*}_{i}\big(\mathbf{G}^{C}\big)=\underbrace{(\alpha-\bar{c})^{2}}_{\text{constant}}\;\underbrace{\left[\phi(n+1)^{2}-1\right]\phi}_{\text{multiplier}}\;\underbrace{\left(\frac{1}{(n+1)^{2}\phi-\sum_{j=3}^{n}\theta_{j}^{2}}\right)^{2}}_{\text{target}},$
$\lim_{\theta_{j}\to 0}\lambda=\infty,\qquad\lim_{\theta_{j}\to 0}\Delta=\frac{4}{(n+1)^{2}\phi-4}.$
Therefore,
$\lim_{\theta_{j}\to 0}{\pi}^{*}_{i}\big(\mathbf{G}^{C}_{-ij}\big)=\underbrace{(\alpha-\bar{c})^{2}}_{\text{constant}}\;\underbrace{\left[\phi\left(\tfrac{n+1}{2}\right)^{2}-1\right]\phi}_{\text{multiplier}}\;\underbrace{\left(\frac{2}{(n+1)^{2}\phi-\sum_{j=3}^{n}\theta_{j}^{2}}\right)^{2}}_{\text{target}}.$
Comparing the two limits yields
$\lim_{\theta_{j}\to 0}{\pi}^{*}_{i}\big(\mathbf{G}^{C}_{-ij}\big)>\lim_{\theta_{j}\to 0}{\pi}^{*}_{i}\big(\mathbf{G}^{C}\big).$
Define
$r(\theta_{j}):=\frac{{\pi}^{*}_{i}\big(\mathbf{G}^{C}\big)}{{\pi}^{*}_{i}\big(\mathbf{G}^{C}_{-ij}\big)}.$
We have shown that
$\lim_{\theta_{j}\to 0}r(\theta_{j})<1\quad\text{and}\quad\lim_{\theta_{j}\to 1}r(\theta_{j})>1.$
By continuity and monotonicity of $r(\theta_{j})$, there exists a threshold $\bar{\theta}_{2}\in(0,1)$ such that $r(\bar{\theta}_{2})=1$. Hence, for $\theta_{j}<\bar{\theta}_{2}$, deviation from the fully connected network is profitable, and the complete network is not pairwise stable.

### v2

@@ Appendix A Proofs
Without loss of generality, fix a productivity distribution $\{\theta_{i}\}_{i\in\mathcal{N}}\in(0,1]^{n}$ such that $1=\theta_{1}\geq\cdots\geq\theta_{n}>0$. Lemma 3 guarantees the existence of a set of thresholds $\{\theta_{ij}^{*}\}_{i\neq j\in\mathcal{N}}$ with $\theta_{ij}^{*}\in(0,\theta_{i})$. On the one hand, notice that for each firm $j\in\mathcal{N}\setminus\{n\}$, all firms $i\in\mathcal{N}$ with $i>j$ would maintain the collaboration link $(i,j)$ since $\theta_{ij}^{*}<\theta_{i}\leq\theta_{j}$. On the other hand, for each firm $j\in\mathcal{N}\setminus\{1\}$, all firms $i\in\mathcal{N}$ with $i<j$ would break the link $(i,j)$ from the complete network if $\theta_{j}<\theta_{ij}^{*}<\theta_{i}$. For each firm $j\in\mathcal{N}\setminus\{1\}$, we define $\Theta_{j}\coloneqq\{\theta_{ij}^{*}\mid i<j\}$ and $\theta_{j}^{*}\coloneqq\max\Theta_{j}$. We then define an economy-wide $\Theta\coloneqq\{\theta_{j}^{*}\mid 1<j\leq n\}$. Note that no firm $i\neq 1$ wishes to break the link to the most productive firm 1 as $\theta_{i1}^{*}<\theta_{i}\leq 1$.
First, let $\theta^{*}\coloneqq\min\Theta$. If there exists any firm $j\in\mathcal{N}\setminus\{1\}$ with productivity level $\theta_{j}<\theta^{*}$, then $\theta_{j}<\theta^{*}\leq\theta_{j}^{*}=\theta_{ij}^{*}$ for some firm $i\in\mathcal{N}$ with $i<j$. Therefore, firm $i$ does not wish to maintain link $(i,j)$, leading to $\mathbf{G}^{C}$ not being pairwise stable. Second, let $\theta^{**}\coloneqq\max\Theta$. If all firms $j\in\mathcal{N}\setminus\{1\}$ satisfy $\theta_{j}\geq\theta^{**}$, then $\theta_{j}\geq\theta_{j}^{*}\geq\theta_{ij}^{*}$ for all firms $i\in\mathcal{N}$ with $i<j$. Therefore, all links are sustainable and the complete network remains pairwise stable. The proof is completed by noting $\theta^{*}\leq\theta^{**}$ as they are the minimum and maximum, respectively, over the set $\Theta$. ∎
Lemma 1.
The complete network $\mathbf{G}^{C}$ is characterized by $G_{ij}^{C}=1$ for all $i,j\in\mathcal{N}$ with $i\neq j$. The effort level of this configuration can be derived in closed form as
${e}^{*}_{i}(\mathbf{G}^{C})=\frac{(\alpha-\bar{c})\theta_{i}}{(n+1)^{2}\phi-\sum_{j=1}^{n}\theta_{j}^{2}}.$ (14)
Proof.
The FOC (6) for the complete network structure can be re-expressed as
${e}^{*}_{i}(\mathbf{G}^{C})$ $=\frac{\theta_{i}}{(n+1)^{2}\phi-\theta_{i}^{2}}\left[\alpha-\bar{c}+\sum_{j\neq i}\theta_{j}{e}^{*}_{j}(\mathbf{G}^{C})\right],$
${e}^{*}_{i}(\mathbf{G}^{C})$ $=\frac{\theta_{i}}{(n+1)^{2}\phi}\left[\alpha-\bar{c}+\sum_{j=1}^{n}\theta_{j}{e}^{*}_{j}(\mathbf{G}^{C})\right],$
where we use the fact that $\eta_{i}(\mathbf{G}^{C})=[n-d_{i}(\mathbf{G}^{C})]/(n+1)=1/(n+1)$ for all $i\in\mathcal{N}$. Pre-multiplying this equation by $\theta_{j}$ and summing across $j\in\mathcal{N}$ leads to
$\sum_{j=1}^{n}\theta_{j}{e}^{*}_{j}(\mathbf{G}^{C})=\frac{(\alpha-\bar{c})\sum_{j=1}^{n}\theta_{j}^{2}}{(n+1)^{2}\phi-\sum_{j=1}^{n}\theta_{j}^{2}}.$
Replacing this result into the FOC and re-arranging terms yields the closed-form solution for complete network efforts (14). ∎
Lemma 2.
Consider a network $\mathbf{G}^{C}_{-kl}$, which removes the link between arbitrary firms $k$ and $l$ while preserving all other connections from the complete network. The effort levels for all firms in this configuration can be derived in closed form as (for $i\in\mathcal{N}\setminus\{k,l\}$)
${e}^{*}_{k}(\mathbf{G}^{C}_{-kl})$ $=\frac{\alpha-\bar{c}}{b_{k}(1-Q)-2\theta_{k}Q+\theta_{l}\lambda[n(1-Q)-(1+Q)]},$ (15)
${e}^{*}_{l}(\mathbf{G}^{C}_{-kl})$ $=\frac{(\alpha-\bar{c})\lambda}{b_{k}(1-Q)-2\theta_{k}Q+\theta_{l}\lambda[n(1-Q)-(1+Q)]},$ (16)
${e}^{*}_{i}(\mathbf{G}^{C}_{-kl})$ $=\frac{(\alpha-\bar{c})\theta_{i}[b_{k}+2\theta_{k}+(n+1)\theta_{l}\lambda]}{(n+1)^{2}\phi\{b_{k}(1-Q)-2\theta_{k}Q+\theta_{l}\lambda[n(1-Q)-(1+Q)]\}},$ (17)
where we define
@@
$b_{k}\coloneqq\frac{(n+1)^{2}\phi}{2\theta_{k}}-2\theta_{k},\quad\lambda\coloneqq\frac{\theta_{l}}{\theta_{k}}\cdot\frac{(n+1)\phi-2\theta_{k}^{2}}{(n+1)\phi-2\theta_{l}^{2}},\quad\text{and}\quad Q\coloneqq\frac{\sum_{j\in\mathcal{N}\setminus\{k,l\}}\theta_{j}^{2}}{(n+1)^{2}\phi}.$
Proof.
Without loss of generality, we set $k=1$ and $l=2$ to consider the network $\mathbf{G}^{C}_{-12}$. In such case, we have $\eta_{1}(\mathbf{G}^{C}_{-12})=\eta_{2}(\mathbf{G}^{C}_{-12})=2/(n+1)$ and $\eta_{i}(\mathbf{G}^{C}_{-12})=1/(n+1)$ for $i\geq 3$. We can then rewrite the system of first-order conditions (6) as
$b_{1}{e}^{*}_{1}(\mathbf{G}^{C}_{-12})$ $=(\alpha-\bar{c})+\sum_{j=3}^{n}\theta_{j}{e}^{*}_{j}(\mathbf{G}^{C}_{-12})-(n-1)\theta_{2}{e}^{*}_{2}(\mathbf{G}^{C}_{-12}),$ (18)
$b_{2}{e}^{*}_{2}(\mathbf{G}^{C}_{-12})$ $=(\alpha-\bar{c})+\sum_{j=3}^{n}\theta_{j}{e}^{*}_{j}(\mathbf{G}^{C}_{-12})-(n-1)\theta_{1}{e}^{*}_{1}(\mathbf{G}^{C}_{-12}),$ (19)
$b_{i}{e}^{*}_{i}(\mathbf{G}^{C}_{-12})$ $=(\alpha-\bar{c})+\sum_{j=3}^{n}\theta_{j}{e}^{*}_{j}(\mathbf{G}^{C}_{-12})+2\theta_{1}{e}^{*}_{1}(\mathbf{G}^{C}_{-12})+2\theta_{2}{e}^{*}_{2}(\mathbf{G}^{C}_{-12}),\quad\text{for }i\geq 3$ (20)
where we define the coefficients $b=(b_{1},\ldots,b_{n})$ as
$b_{1}=\frac{(n+1)^{2}\phi}{2\theta_{1}}-2\theta_{1},\quad b_{2}=\frac{(n+1)^{2}\phi}{2\theta_{2}}-2\theta_{2},\quad\text{and}\quad b_{i}=\frac{(n+1)^{2}\phi}{\theta_{i}},\quad\text{for }i\geq 3.$
Equating the right-hand sides of (18) to (19), we can obtain
${e}^{*}_{2}(\mathbf{G}^{C}_{-12})=\frac{b_{1}-(n-1)\theta_{1}}{b_{2}-(n-1)\theta_{2}}{e}^{*}_{1}(\mathbf{G}^{C}_{-12})\implies{e}^{*}_{2}(\mathbf{G}^{C}_{-12})=\lambda{e}^{*}_{1}(\mathbf{G}^{C}_{-12}),$ (21)
where we define $\lambda\coloneqq[b_{1}-(n-1)\theta_{1}]/[b_{2}-(n-1)\theta_{2}]=(\theta_{2}/\theta_{1})\cdot[(n+1)\phi-2\theta_{1}^{2}]/[(n+1)\phi-2\theta_{2}^{2}]$ to represent this ratio. Now, equate the right-hand sides of (18) to (20) and follow similar steps to obtain
${e}^{*}_{i}(\mathbf{G}^{C}_{-12})=\frac{b_{1}+2\theta_{1}+(n+1)\theta_{2}\lambda}{b_{i}}{e}^{*}_{1}(\mathbf{G}^{C}_{-12}),\quad\text{for }i\geq 3.$ (22)
Define $T$ as the sum over effective efforts for unaffected firms, such that $T\coloneqq\sum_{j=3}^{n}\theta_{j}{e}^{*}_{j}(\mathbf{G}^{C}_{-12})$. From (22) we can obtain
$T$ $=[b_{1}+2\theta_{1}+(n+1)\theta_{2}\lambda]\cdot Q\cdot{e}^{*}_{1}(\mathbf{G}^{C}_{-12}).$ (23)
where we define $Q\coloneqq\sum_{j=3}^{n}\theta_{j}/b_{j}=\sum_{j=3}^{n}\theta_{j}^{2}/[(n+1)^{2}\phi]$. Replace (21) and (23) back into (18) to obtain an expression for the effort of firm 1:
$b_{1}{e}^{*}_{1}(\mathbf{G}^{C}_{-12})$ $=(\alpha-\bar{c})+T-(n-1)\theta_{2}\lambda{e}^{*}_{1}(\mathbf{G}^{C}_{-12}),$
${e}^{*}_{1}(\mathbf{G}^{C}_{-12})$ $=\frac{\alpha-\bar{c}}{b_{1}(1-Q)-2\theta_{1}Q+\theta_{2}\lambda[n(1-Q)-(1+Q)]}.$ (24)
Plugging (24) into (21) and (22), we can obtain the efforts for all remaining firms: ${e}^{*}_{2}(\mathbf{G}^{C}_{-12})$ as in (16) and ${e}^{*}_{i}(\mathbf{G}^{C}_{-12})$ for $i\geq 3$ as in (17). With some algebra, these solutions can be expressed in terms of fundamental model quantities as (for $i\geq 3$)
${e}^{*}_{1}(\mathbf{G}^{C}_{-12})$ $=\frac{2(\alpha-\bar{c})\theta_{1}[(n+1)\phi-2\theta_{2}^{2}]}{(n+1)\phi[(n+1)^{2}\phi(1-Q)-4(\theta_{1}^{2}+\theta_{2}^{2})]+4\theta_{1}^{2}\theta_{2}^{2}[(n+1)Q-n+3]},$
${e}^{*}_{2}(\mathbf{G}^{C}_{-12})$ $=\frac{2(\alpha-\bar{c})\theta_{2}[(n+1)\phi-2\theta_{1}^{2}]}{(n+1)\phi[(n+1)^{2}\phi(1-Q)-4(\theta_{1}^{2}+\theta_{2}^{2})]+4\theta_{1}^{2}\theta_{2}^{2}[(n+1)Q-n+3]},$
${e}^{*}_{i}(\mathbf{G}^{C}_{-12})$ $=\frac{(\alpha-\bar{c})\theta_{i}[(n+1)^{2}\phi^{2}-4(\theta_{1}^{2}+\theta_{2}^{2})]}{(n+1)^{2}\phi^{2}[(n+1)^{2}\phi(1-Q)-4(\theta_{1}^{2}+\theta_{2}^{2})]+4(n+1)\phi\theta_{1}^{2}\theta_{2}^{2}[(n+1)Q-n+3]}.$
Lemma 3 (Link deviation in the complete network).
For any pair of firms $(i,j)\in\mathcal{N}$ in a complete network $\mathbf{G}^{C}$, there exists a threshold over firm $j$’s productivity $\theta_{j}$, denoted by ${\theta}^{*}_{ij}(\theta_{1},\ldots,\theta_{j-1},\theta_{j+1},\ldots,\theta_{n})\in(0,\theta_{i})$, such that:
$\pi^{*}_{i}\big(\mathbf{G}^{C}_{-ij}\big)>\pi^{*}_{i}\big(\mathbf{G}^{C}\big)\iff\theta_{j}<{\theta}^{*}_{ij}(\theta_{1},\ldots,\theta_{j-1},\theta_{j+1},\ldots,\theta_{n}).$ (25)
Proof of Lemma 3.
Without loss of generality, consider again the situation where firm 1 is in consideration for sustaining or breaking a link with firm 2. Using the equilibrium profit expression (9) in combination with Lemmas 1 and 2, the profit for the firm 1 under networks $\mathbf{G}^{C}$ and $\mathbf{G}_{-12}^{C}$ is given by
${\pi}^{*}_{1}(\mathbf{G}^{C})$ $=\left[\frac{\phi}{\theta_{1}^{2}}(n+1)^{2}-1\right]\phi\left[\frac{(\alpha-\bar{c})\theta_{1}}{(n+1)^{2}\phi-\sum_{j=1}^{n}\theta_{j}^{2}}\right]^{2},$
${\pi}^{*}_{1}(\mathbf{G}^{C}_{-12})$ $=\left[\frac{\phi}{\theta_{1}^{2}}\left(\frac{n+1}{2}\right)^{2}-1\right]\phi\left\{\frac{\alpha-\bar{c}}{b_{1}(1-Q)-2\theta_{1}Q+\theta_{2}\lambda[n(1-Q)-(1+Q)]}\right\}^{2}.$
Define $R\coloneqq{\pi}^{*}_{1}(\mathbf{G}^{C}_{-12})/{\pi}^{*}_{1}(\mathbf{G}^{C})$ as ratio of profits for firm 1 between of the network with a deviation to the complete network. Using the definitions of $b_{1}$ and $Q$, we can then express
$R=\frac{(n+1)^{2}\phi-4\theta_{1}^{2}}{(n+1)^{2}\phi-\theta_{1}^{2}}\left\{\frac{(n+1)^{2}\phi(1-Q)-\theta_{1}^{2}-\theta_{2}^{2}}{(n+1)^{2}\phi(1-Q)-4\theta_{1}^{2}+2\theta_{1}\theta_{2}\lambda[(n+1)(1-Q)-2]}\right\}^{2}.$
First, we show that $\lim_{\theta_{2}\to 0}R>1$, proving that there exists a profitable deviation from complete when $\theta_{2}\approx 0$ regardless of the remaining productivity distribution $\{\theta_{j}\}_{j=3}^{n}$. With $\theta_{2}$ approaching 0, the ratio simplifies to
$\lim_{\theta_{2}\to 0}R$ $=\frac{(n+1)^{2}\phi-4\theta_{1}^{2}}{(n+1)^{2}\phi-\theta_{1}^{2}}\left\{\frac{(n+1)^{2}\phi(1-Q)-\theta_{1}^{2}}{(n+1)^{2}\phi(1-Q)-4\theta_{1}^{2}}\right\}^{2},$
$=\underbrace{\frac{(n+1)^{2}\phi-4\theta_{1}^{2}}{(n+1)^{2}\phi(1-Q)-4\theta_{1}^{2}}\cdot\frac{(n+1)^{2}\phi(1-Q)-\theta_{1}^{2}}{(n+1)^{2}\phi-\theta_{1}^{2}}}_{=g(4)/g(1)}\cdot\underbrace{\frac{(n+1)^{2}\phi(1-Q)-\theta_{1}^{2}}{(n+1)^{2}\phi(1-Q)-4\theta_{1}^{2}}}_{>1},$
where we define $g(x):=[(n+1)^{2}\phi-x\cdot\theta_{1}^{2}]/[(n+1)^{2}\phi(1-Q)-x\cdot\theta_{1}^{2}]$. We next show that $g(x)$ is a strictly increasing function, which makes the product of the first two terms $g(4)/g(1)>1$, guaranteeing $\lim_{\theta_{2}\to 0}R>1$. Taking the derivative of $g$ with respect to $x$, and using $n\geq 3$, $\phi\geq 1$, $x\in[1,4]$, and $Q>0$, we have
$\frac{dg(x)}{dg}=\frac{\theta_{1}^{2}(n+1)^{2}\phi Q}{[(n+1)^{2}\phi-x]^{2}}>0.$
Second, we will show $\lim_{\theta_{2}\to\theta_{1}}R<1$, such that that the complete network is pairwise stable when $\theta_{2}\approx\theta_{1}$ regardless of the remaining productivity distribution $\{\theta_{j}\}_{j=3}^{n}$. With $\theta_{2}$ approaching $\theta_{1}$, $\lambda$ approaches 1 and the ratio $R$ simplifies to
$\lim_{\theta_{2}\to\theta_{1}}R$ $=\underbrace{\frac{(n+1)^{2}\phi-4\theta_{1}^{2}}{(n+1)^{2}\phi-\theta_{1}^{2}}}_{<1}\Bigg\{\underbrace{\frac{(n+1)^{2}\phi(1-Q)-2\theta_{1}^{2}}{(n+1)^{2}\phi(1-Q)-2\theta_{1}^{2}+2\theta_{1}^{2}[(n+1)(1-Q)-3]}}_{\coloneqq\rho}\Bigg\}^{2}.$
We are left with showing $\rho<1$. As $\theta_{i}\leq 1$ for all $i\in\mathcal{N}$, $Q\leq(n-2)/[(n+1)^{2}\phi]$, which implies $(n+1)^{2}\phi(1-Q)\geq(n+1)^{2}\phi-n+2\geq 2\geq 2\theta_{1}^{2}$. It is then sufficient to show (n+1)(1Q)>3(n+1)(1-Q)>3 to guarantee $\rho<1$. Note that we have
$(n+1)(1-Q)\geq n+1-\frac{n-2}{(n+1)\phi}>n\geq 3.$
Finally, note that the bounds introduced on fundamental model quantities ensure the denominator of $R$ is bounded away from 0. As a rational function of $\theta_{2}$, $R$ is therefore continuous on $\theta_{2}$, with $\lim_{\theta_{2}\to 0}R>1$ and $\lim_{\theta_{2}\to\theta_{1}}R<1$. The derivative of $R$ can be signed using the same bounds as in previous inequalities:
@@
$\frac{\partial R}{\partial\theta_{2}}$ $=2\cdot\frac{(n+1)^{2}\phi-4\theta_{1}^{2}}{(n+1)^{2}\phi-\theta_{1}^{2}}\cdot\frac{N}{D^{3}}\left(\frac{\partial N}{\partial\theta_{2}}\cdot D-N\cdot\frac{\partial D}{\partial\theta_{2}}\right),\quad\text{with}$
$\frac{\partial N}{\partial\theta_{2}}\cdot D-N\cdot\frac{\partial D}{\partial\theta_{2}}$ $=-2\left\{\theta_{2}D+N[(n+1)(1-Q)-2]\frac{2(n+1)\phi\lambda}{(n+1)\phi-2\theta_{2}^{2}}\right\}<0,$
where $N\coloneqq(n+1)^{2}\phi(1-Q)-\theta_{1}^{2}-\theta_{2}^{2}$, and $D\coloneqq(n+1)^{2}\phi(1-Q)-4\theta_{1}^{2}+2\theta_{1}\theta_{2}\lambda[(n+1)(1-Q)-2]$ are the numerator and denominator, respectively, in the relative efforts squared term of $R$. Monotonicity and the intermediate value theorem guarantee the existence of a unique threshold $0<{\theta}^{*}_{12}<1$ where $R=1$, at which the most productive firm 1 is indifferent between keeping or removing the R&D collaboration to firm 2. That is, the complete network is not pairwise stable for all $\theta_{2}<{\theta}^{*}_{12}$, while 1 wants to maintain the link to 2 if $\theta_{2}\geq{\theta}^{*}_{12}$. ∎